Calculate consumption of fixed capital
NNPfc 6000 GDPmp 8000 NIT 400 NFIA 100
1
Expert's answer
2020-08-28T10:49:03-0400

Сonsumption of fixed capital= Depreciation


NDPFC = GDPMP – Depreciation – Net Indirect Taxes


National Income (NNPFC) = NDPFC + NFIA


NDPFC=National Income (NNPFC) -NFIA=6000-100=5900


5900=8000- Depreciation-400


Depreciation=8000-5900-400=1700